# Learning Knowledge-based Prompts for Robust 3D Mask Presentation Attack Detection

### v1

@@ Abstract
Face recognition systems, known for their high accuracy and convenience, have been deeply integrated into everyday applications, including residential access control, attendance systems, and criminal tracking. However, face presentation attacks pose a significant challenge to the reliability of these systems [1, 2]. Imposers commonly exploit printed photos, replayed videos, or 3D masks to impersonate real faces and deceive face recognition systems. To counter these threats, the task of face presentation attack detection (PAD), which aims at distinguishing fake faces from real ones, has drawn considerable attention from both academia and industry [3, 4].
@@ 2 Related Work
2.1.1 Single Modality-based Detection Methods
@@ 2.2 Vision-Language Models for Attack Detection
In light of recent advances in natural language processing [43], numerous methods focus on text prompt learning. For example, CoOp [21] introduces learnable word vectors to capture context information for various classes. CoCoOp [22] builds upon CoOp, generating image-specific text prompts through a meta-learning approach to reduce overfitting on known classes. Unlike text prompt learning, VPL [44] involves learning an image perturbation added to input images as visual prompts. Additionally, unified prompt tuning [45] and multi-modal prompt learning [46] integrate text and visual prompt learning simultaneously, combining both benefits. Recently, prompt learning has been applied to advanced vision task face presentation attack detection [23, 47, 24, 48, 49].
@@
Srivatsan et al. [23] design a collection of fixed category prompts for real and fake faces and fine-tune all parameters of the CLIP model to align image and text features for face presentation attack detection. Yu et al. [47] learn the modal-relevant visual prompts to explore the ability of the frozen foundation model for the problems of partial modal absence in multimodality-based face presentation attack detection. Meanwhile, textually prompt learning has been introduced for domain generalization-based face presentation attack detection in [24, 48, 49]. Fang et al. [24] provide fine-grained textual descriptions of the face region to reduce the model’s focus on irrelevant facial information and develop sample-level image-text alignment strategies to learn domain-invariant features. Different semantic prompts conditioned on content and style features are learned by two lightweight transformers in [48]. Additionally, Mu et al. [49] construct matching and non-matching textual prompts and use these cross-domain generic textual prompts to explicitly constrain the model focusing on domain-invariant features rather than domain-specific details such as lighting conditions, and recording devices.
@@
In contrast to existing prompt learning-based face presentation attack detection methods, which concentrate on detecting 2D spoof faces, we focus on detecting high-fidelity 3D masks. These masks are more similar to real faces in color, texture, and three-dimensional structure and are thus more difficult to detect. Moreover, we inject category-related knowledge graphs into prompt learning, while constructing a spurious correlation elimination learning paradigm to jointly mine the pre-trained knowledge of the foundation model adapted to the detection task.
@@ 3 Proposed Method
Denote the available training domain as , where represents the -th data sample, and is its corresponding label, belonging to the label set , and denotes the total number of classes. The label set includes two classes: real faces and 3D masks. Our objective is to learn a 3D mask presentation attack detection model trained on , with the capability to generalize effectively to an unseen testing domain . The testing domain shares the same label set with the training domain but may exhibit a different data distribution due to variations in factors such as recording scenes, devices, lighting conditions, and the appearance or material of 3D masks.
@@ 3.1 Knowledge Graph Prompt Scheme
Common knowledge graphs, such as ConceptNet and Wikidata, encapsulate a wealth of expert domain knowledge. This knowledge is typically a comprehensive and well-organized summary of experts from a large collection of samples. For instance, mask materials may include paper, plastic, silicone, and resin. Such extensive and structured knowledge undoubtedly aids in learning generalized features for 3D mask presentation attack detection. We design a knowledge graph prompt tailored for the vision-language model to leverage this.
MaskPAD-KG construction. We first use class names in to query the common knowledge graphs such as ConcepNet and Wikidata, retrieving category-related entities and relationships. To ensure the accuracy of the resulting knowledge graph, we manually review the retrieved entities and relationships. While most common knowledge graphs are constructed based on general knowledge, they may lack task-specific information critical for 3D mask presentation attack detection. To address this, we enhance the retrieved subgraphs by adding supplementary entities based on common clues in 3D mask presentation attack detection, focusing on three key aspects: category-related terms, symbolic meanings, and inherent characteristics. Overall, the workload for knowledge graph construction remains minimal, as we only need to create graphs for the two classes: real faces and 3D masks.
The constructed knowledge graph for 3D mask presentation attack detection (MaskPAD-KG) is illustrated in Figure 3. MaskPAD-KG is represented as , where denotes the set of entities in the knowledge graph, represents the set of relationships, and represents the knowledge in the form of triples . Here, and are the head and tail entities, respectively, and is their relationship. Both the entity set and the triple set encapsulate domain-specific knowledge essential for 3D mask presentation attack detection. We encode this knowledge into prompts by leveraging both the entity set and the knowledge triples.
Knowledge graph prompt generation. Given the -th class name in (), we combine all category-related entities in following the pattern shown in Equation 1 to generate the category-associated entity prompt .
@@
Large language models (e.g., GPT-4) trained on vast datasets contain a wealth of knowledge. Given the -th class name in (), we extract the knowledge from the large language model to automatically generate the description for every triple related to . The question template used to query the large language model is as follows: “In 3D mask face presentation attack detection, please analyze the sentence [][][] in 30 words or less.” One example is illustrated in Figure 4. The fine-grained discriminative description for is formally defined as:
where represents the description corresponding to the -th triple in associated with the -th class name , is the number of descriptions associated with the -th class name.
@@
Visual-specific knowledge filter. A single input image is typically linked to only a subset of the information in the knowledge graph. The focal points in the mapping between different images and the knowledge graph prompts (entities or fine-grained discriminative descriptions) may differ. To better align the image and text modalities, we further refine the prompts into visual-specific knowledge graph prompts conditioned on the image content. A visual-specific knowledge filter is designed to dynamically assign varying weights to the entities or the fine-grained discriminative descriptions based on their relevance to the images, with higher weights assigned to more relevant ones.
Specifically, to better adapt image features to the 3D mask presentation attack detection task and align with the text prompts, we employ an adaptation layer, which can be implemented simply with two fully connected layers, to transform the image features into the output representation . Let be the token embedding of the knowledge graph prompt (i.e., the entity prompt or fine-grained discriminative description). is the dimension of the last dimension of . The linear projection layer is denoted as . For a prompt , the calculation of the visual-specific prompt by the visual-specific knowledge filter is as follows:
Considering the potential differences in alignment between knowledge graph prompts and images, as illustrated in Figure 2, we design separate visual-specific knowledge filters for the entities and the fine-grained discriminative descriptions, denoted as and , respectively. The visual-specific knowledge graph prompt for the -th class name is formed by concatenating the refined entity and the fine-grained discriminative description.
The generated visual-specific knowledge prompt can comprehensively utilize knowledge in the knowledge graph while selectively leveraging this knowledge to provide accurate contextual information automatically. This approach facilitates mining useful pre-trained knowledge from vision-language models for 3D mask presentation attack detection.
In addition to the visual-specific knowledge graph prompt, we also design the learnable context prompt. Given the -th class name in (), its prompt consists of the token embedding of the prefix context sequence and the class name:
where () is the -th context vector and is the number of context tokens. Ultimately, the visual-specific knowledge graph prompt and the learnable context prompt are concatenated and fed into the text encoder to obtain the knowledge-based text feature .
@@ 3.2 Spurious Correlation Elimination Prompt Learning
Causal graph analysis. Figure 5(a) illustrates the key elements influencing 3D mask presentation attack detection. Here, nodes and represent the observable input image and its label, while nodes and denote unobservable factors relevant and irrelevant to 3D mask presentation attack detection. The detection-relevant factors capture critical attributes that differentiate real faces from 3D masks, such as material and facial structure. These attributes typically exhibit strong cross-scenario generalization capabilities. In contrast, the detection-irrelevant factors represent external, scenario-specific influences, such as the recording environment and equipment. While these factors do not pertain to the classification of real faces versus 3D masks, they significantly affect the appearance of the image and vary across different scenarios. Both and jointly shape the observable presentation of the image .
@@
The label should be determined exclusively by the detection-relevant factors . However, since is unobservable, models must rely on the relationship between and . This reliance creates vulnerability to domain shifts, as the widely used empirical risk minimization objective inherently conditions on . Consequently, spurious correlations emerge between and through the causal pathway . As a result, the learned models often incorporate detection-irrelevant yet environment-specific information, impairing the generalization capability. To address this issue, it is essential to eliminate the correlation between and , as illustrated in Figure 5(b). We propose enforcing a causal independence property, , which ensures that is independent of . This approach constrains the model to focus solely on detection-relevant features, excluding environment-specific information. By doing so, we enhance the model’s robustness and generalization capability.
@@
Spurious correlation elimination regularization. In this work, we represent the label as category-related prompts. To enforce the causal independence property , we propose a spurious correlation elimination regularization that constrains the generated prompts, requiring them to be independent of detection-irrelevant factors . Previous works have shown that local detailed information can provide effective discriminative information for 3D mask presentation attack detection [50, 51]. As such, our approach mitigates spurious correlations between the generated prompts and category-irrelevant local patches within the image.
To determine the relevance of each local patch, we calculate its similarity with the prompts of all categories. The similarity between patch and prompt related to category is computed as follows:
@@
where represents the local feature of the patch , and represents the knowledge-based text feature of category , and is the temperature coefficient.
@@
A patch is deemed category-irrelevant if its most similar category does not match the ground truth label . The set of category-irrelevant patches for a sample is defined as:
The spurious correlation elimination regularization is designed to maximize the entropy of the similarity :
Note that represents the probability that patch belongs to the category . Maximizing the entropy encourages a uniform probability distribution of over the label set, reducing the likelihood that category-irrelevant patches align with prompts corresponding to the ground truth. This ensures that the generated prompts are decoupled from category-irrelevant patches, effectively mitigating spurious correlations, minimizing the impact of detection-irrelevant factors.
@@
Spoof-relevant discriminant regularization. In addition to the spurious correlation elimination regularization, we design a spoof-relevant discriminant regularization to align the image features with the corresponding knowledge-based text features of the ground truth. Formally, is implemented as the cross-entropy loss function:
where
@@
and denotes a binary indicator, if is the -th class of , and otherwise. The spoof-relevant discriminant regularization encourages the model to assign the highest predictive probability to the class corresponding to the ground truth, effectively aligning image features with text features. It guides the model to focus on detection-relevant factors, enhancing robustness across diverse scenarios.

### v2

@@ Abstract
Face recognition systems, known for their high accuracy and convenience, have been deeply integrated into everyday applications, including residential access control, attendance systems, and criminal tracking. However, face presentation attacks pose a significant challenge to the reliability of these systems [1]. Imposers commonly exploit printed photos, replayed videos, or 3D masks to impersonate real faces and deceive face recognition systems. To counter these threats, the task of face presentation attack detection (PAD), which aims at distinguishing fake faces from real ones, has drawn considerable attention from both academia and industry [2, 3].
@@ 2 Related Work
In this section, we review related work from three perspectives: 3D mask presentation attack detection, vision-language model-based attack detection, and evaluation metrics commonly used for face presentation attack detection.
@@
Based on the modality of the input image, we categorize 3D mask presentation attack detection methods into two main groups: single-modality and multimodality-based approaches.
2.1.1 Single-Modality-based Detection Methods
@@ 2.2 Vision-Language Models for Attack Detection
In light of recent advances in natural language processing [42], numerous methods focus on text prompt learning. For example, CoOp [20] introduces learnable word vectors to capture context information for various classes. CoCoOp [21] builds upon CoOp, generating image-specific text prompts through a meta-learning approach to reduce overfitting on known classes. Unlike text prompt learning, VPL [43] involves learning an image perturbation added to input images as visual prompts. Additionally, unified prompt tuning [44] and multi-modal prompt learning [45] integrate text and visual prompt learning simultaneously, combining both benefits. Recently, prompt learning has been applied to advanced vision task face presentation attack detection [22, 46, 23, 47, 48, 49].
@@
Srivatsan et al. [22] design a collection of fixed category prompts for real and fake faces and fine-tune all parameters of the CLIP model to align image and text features for face presentation attack detection. Yu et al. [46] learn the modal-relevant visual prompts to explore the ability of the frozen foundation model for the problems of partial modal absence in multimodality-based face presentation attack detection. Meanwhile, textually prompt learning has been introduced for domain generalization-based face presentation attack detection in [23, 47, 49, 48]. Fang et al. [23] provide fine-grained textual descriptions of the face region to reduce the model’s focus on irrelevant facial information and develop sample-level image-text alignment strategies to learn domain-invariant features. Different semantic prompts conditioned on content and style features are learned by two lightweight transformers in [47]. Similarly, the learnable prompt tokens are required to carry diverse visual styles from implicit synthesis of mixed novel styles in [48], enhancing the generalizability of the learned prompts to domain-specific stylistic variations. Additionally, Mu et al. [49] construct matching and non-matching textual prompts and use these cross-domain generic textual prompts to explicitly constrain the model focusing on domain-invariant features rather than domain-specific details such as lighting conditions, and recording devices.
@@
In contrast to existing prompt learning-based face presentation attack detection approaches that primarily focus on detecting 2D spoof faces, our method targets the more challenging task of high-fidelity 3D mask detection. These masks exhibit high similarity to real faces in terms of color, texture, and structure, making them significantly more difficult to distinguish. To address this challenge, we inject explicit and interpretable knowledge graphs into prompt learning, while constructing a spurious correlation elimination learning paradigm to jointly adapt general knowledge of foundation models for the detection task.
2.3 Evaluation Metrics for Face Presentation Attack Detection
The performance evaluation of face presentation attack detection algorithms requires a comprehensive assessment of error rates associated with both real faces and presentation attacks. The evaluation metrics commonly used in this field can be broadly categorized into two principal groups:
The first category comprises traditional biometric classification metrics [50], including False Acceptance Rate (FAR), False Rejection Rate (FRR), Equal Error Rate (EER), Half Total Error Rate (HTER), and Area Under the Receiver Operating Characteristic Curve (AUC). FAR refers to the proportion of presentation attacks incorrectly accepted as real, while FRR denotes the proportion of real faces incorrectly rejected as presentation attacks. By varying the decision threshold, pairs of FAR and FRR values are obtained, forming the Receiver Operating Characteristic (ROC) curve. EER is calculated on the development set and corresponds to the point on the ROC curve where FAR equals FRR. It reflects the threshold at which the false acceptance and false rejection rates are balanced. HTER is computed on the test set using the decision threshold determined on the development set (i.e., at EER), and is defined as the average of FAR and FRR on the test set. AUC quantifies the area under the ROC curve and serves as a comprehensive measure of overall classification performance.
The second category includes standardized metrics proposed by ISO/IEC 30107-1 [51], namely: Bonafide Presentation Classification Error Rate (BPCER), Attack Presentation Classification Error Rate (APCER), and Average Classification Error Rate (ACER). BPCER captures the error rate associated with misclassifying real presentations. In contrast to FAR, which aggregates all attack types into a single category, APCER evaluates each attack type individually and reports the highest error rate among them, thus reflecting the vulnerability of models to the most challenging attack type. ACER is defined as the average of APCER and BPCER, providing a balanced evaluation that considers both false acceptances and false rejections equally.
@@ 3 Proposed Method
Denote the available training domain as , where represents the -th data sample, and is its corresponding label, belonging to the label set , and denotes the total number of classes. The label set includes two classes: real faces and 3D masks. Our objective is to learn a 3D mask presentation attack detection model trained on , with the capability to generalize effectively to an unseen testing domain . The testing domain shares the same label set with the training domain but may exhibit a different data distribution due to variations in factors such as recording scenes, devices, lighting conditions, and the appearance or material of 3D masks. To achieve this objective, we propose a knowledge-based prompt learning framework, as illustrated in Figure 2.
@@ 3.1 Knowledge Graph Prompt Scheme
Common knowledge graphs, such as ConceptNet and Wikidata [52], encapsulate a wealth of domain expert knowledge. This knowledge is typically a comprehensive and well-organized summary of experts from a large collection of samples. For instance, mask materials may include paper, plastic, silicone, and resin. Such extensive and structured knowledge undoubtedly aids in learning generalized features for 3D mask presentation attack detection. We design a knowledge graph prompt tailored for the vision-language model to leverage this.
MaskPAD-KG construction. We begin by using category names in the label set to query common public knowledge graphs such as ConceptNet and Wikidata, retrieving category-relevant entities and relationships. To ensure the task relevance and accuracy of the resulting subgraph, we manually inspect the retrieved results and remove entities or relationships unrelated to the 3D mask presentation attack detection task (e.g., the entity batman associated with 3D mask). In this work, prior knowledge is organized in three aspects: category-related terms, which frequently co-occur with the category name in contextual descriptions (e.g., pretense, concealment); symbolic meanings, which represent abstract meanings or synonyms assigned to the category by humans in the context of 3D mask presentation attack detection (e.g., unreal, unnatural); and inherent characteristics, which describe the essential properties of the category (e.g., uniform texture, silicone material). These three dimensions allow the knowledge graph to extend beyond surface-level correlations and capture deeper semantic attributes of each category, thereby enriching the category names with more meaningful context.
While most general-purpose knowledge graphs focus on broad commonsense knowledge, they may lack some task-specific details critical for 3D mask presentation attack detection. To address this, we augment the retrieved subgraph by incorporating additional entities based on domain-specific cues commonly observed in 3D mask presentation attack detection. The augmented entities are similarly organized into the aforementioned three semantic dimensions. For example, for the 3D mask category, we add symbolic term spoof face, and inherent characteristics including edge, seam, subtle contour, and rigid shape. Overall, the effort required to construct the knowledge graph is minimal, as it is only performed for two categories: real face and 3D mask.
The final knowledge graph constructed for 3D mask presentation attack detection, referred to as MaskPAD-KG, is illustrated in Figure 3. Formally, MaskPAD-KG is defined as , where denotes the set of entities, and represents the set of relationships. Entities and relationships are combined to form the set of triples . Each triple consists of a head entity , a tail entity , and a relationship linking them. As illustrated in Figure 3, MaskPAD-KG contains a total of 44 entities, 4 types of relationships, and 42 triples. Both the entity set and the triple set encapsulate domain-specific knowledge essential for 3D mask presentation attack detection. We encode this structured knowledge into prompts by leveraging both the entity set and the triple set.
Knowledge graph prompt generation. Given the -th category name in (), we combine all category-related entities in following the pattern shown in Equation 1 to generate the category-associated entity prompt .
@@
Large language models (e.g., GPT-4) trained on vast datasets contain a wealth of knowledge. Given the -th category name in (), we extract the knowledge from the large language model to automatically generate the description for every triple related to . The question template used to query the large language model is as follows: “In 3D mask face presentation attack detection, please analyze the sentence [][][] in 30 words or less.” One example is illustrated in Figure 4. The fine-grained discriminative description for is formally defined as:
where represents the description corresponding to the -th triple in associated with the -th category name , is the number of descriptions associated with the -th category name.
@@
Visual-specific knowledge filter. A single input image is typically associated with only a subset of the information contained in the knowledge graph. The relevance between different images and knowledge graph entities varies, and similar variations exist for the corresponding fine-grained discriminative descriptions. To achieve better alignment between visual and textual modalities, we introduce visual-specific knowledge filters, which dynamically select and weight relevant knowledge terms (both entities and descriptions) by learning a soft attention distribution over candidate prompts, conditioned on the visual content of the input image.
Given the distinct alignment patterns between images and entities, as well as between images and descriptions, as illustrated in Figure 2, we design two separate filtering modules: a visual-specific entity filter and a visual-specific description filter . Each is responsible for selectively weighting its corresponding textual components based on their relevance to the input image. The visual-specific entity filter assigns higher weights to entities that are more semantically aligned with the visual input, while the visual-specific description filter performs a similar function for fine-grained discriminative descriptions. Although these two filters operate on different types of knowledge graph features, they share an identical internal architecture, as shown in Figure 5.
To better adapt the image features for the 3D mask presentation attack detection task and enhance alignment with textual prompts, we employ an adaptation layer composed of two fully connected layers. This module transforms the original visual representation into a task-specific embedding, denoted as . Let represent the token embedding of a textual prompt (either for entities or for descriptions), with denoting the dimensionality of its last dimension. A linear projection layer, denoted by , is then applied to compute the output of a visual-specific knowledge filter for prompt as follows:
For a given category name , we concatenate the entity prompt refined by and the description refined by to form its final visual-specific knowledge graph prompt :
The proposed visual-specific knowledge filters play a critical role in suppressing irrelevant or potentially misleading information that may exist in the raw knowledge graph or generated descriptions. By tailoring the selected knowledge to the visual input, the generated visual-specific knowledge graph prompt offers accurate and context-aware guidance, thereby enhancing both the robustness and generalization ability for 3D mask presentation attack detection.
In addition to the visual-specific knowledge graph prompt, we also design the learnable context prompt. Given the -th category name in (), its prompt consists of the token embedding of the prefix context sequence and the category name:
where () is the -th context vector for and is the number of context tokens. Ultimately, the visual-specific knowledge graph prompt and the learnable context prompt are concatenated and fed into the text encoder to obtain the knowledge-based text feature .
@@ 3.2 Spurious Correlation Elimination Prompt Learning
Causal graph analysis. To provide a principled foundation for the constraint design, we construct a causal graph that explicitly models the dependency structure among critical variables influencing 3D mask presentation attack detection, as illustrated in Figure 6(a). Here, nodes and represent the observable input image and its corresponding label, respectively; represents external confounders, while nodes and refer to unobservable factors relevant and irrelevant to 3D mask presentation attack detection. The detection-relevant factors capture critical attributes that differentiate real faces from 3D masks, such as material and facial structure. These attributes typically exhibit strong cross-scenario generalization capabilities. In contrast, the detection-irrelevant factors represent external, scenario-specific influences, such as the recording environment and equipment. While detection-irrelevant factors do not pertain to the classification of real faces versus 3D masks, they significantly affect the appearance of the image and vary across different scenarios. Both and jointly shape the observable presentation of the image .
@@
The label should be determined exclusively by the detection-relevant factors . However, since is unobservable, models must rely on the relationship between and . This reliance creates vulnerability to domain shifts, as the widely used empirical risk minimization objective inherently conditions on . Consequently, spurious correlations emerge between and through the causal pathway . As a result, the learned models often incorporate detection-irrelevant yet environment-specific information, impairing the generalization capability. To address this issue, it is essential to break the spurious causal path and eliminate the correlation between and , as illustrated in Figure 6(b). We propose enforcing a causal independence property, , which ensures that is independent of . This approach constrains the model to focus solely on detection-relevant features, excluding environment-specific information. By doing so, we enhance the robustness and generalization capability of models.
@@
Spurious correlation elimination regularization. In this work, we represent the label using category-related prompts. To enforce the causal independence property , we introduce a spurious correlation elimination regularization that encourages the generated prompts to be independent of detection-irrelevant factors . An overview of this regularization strategy is presented in Figure 7. Prior studies have shown that local fine-grained visual information can offer effective discriminative cues for 3D mask presentation attack detection [53, 54]. Accordingly, our approach aims to suppress spurious correlations between generated prompts and category-irrelevant local patches within the input image.
To this end, we first divide each input image into non-overlapping local patches. To assess the relevance of each patch for a sample , we compute its similarity with the prompts of all categories. The similarity between the -th patch and the prompt of the -th category is defined as:
@@
where is the local feature of the -th patch, denotes the knowledge-based text feature of the -th category, and is the temperature coefficient.
@@
A patch is deemed category-irrelevant if its most similar category does not correspond to the ground-truth label . Thus, the set of category-irrelevant patches for the sample is defined as:
To mitigate spurious correlations between the generated prompts and the identified category-irrelevant patches, we maximize the entropy of their similarity distribution via the following spurious correlation elimination regularization:
Here, can be interpreted as the probability of the -th patch belonging to the -th category. Maximizing the entropy enforces a uniform distribution across categories, thereby reducing the chance that generated prompts spuriously align with irrelevant patches. This promotes the independence of generated prompts from category-irrelevant content and mitigates the influence of detection-irrelevant factors.
@@
Spoof-relevant discriminant regularization. Beyond eliminating spurious correlations, it is equally important to strengthen the correlations between detection-relevant factors and ground-truth labels. To this end, we introduce a spoof-relevant discriminant regularization , which aligns image features with the knowledge-based text features of the ground-truth category. Formally, is defined as a cross-entropy loss:
Here, is a binary indicator such that if is the -th category in the label set , and otherwise. is the probability assigning the input image to the -th category, which is computed as:
The value of reflects the similarity between the image feature of the input and the text feature associated with the -th category, and can be interpreted as the predicted probability for that category.
@@
This regularization encourages the model to produce a peaked probability distribution over the categories, with the highest probability assigned to the ground-truth class. It promotes maximum alignment between the image features and the text features of the corresponding category, thus reinforcing the model’s focus on detection-relevant factors. Consequently, the spoof-relevant discriminant regularization enhances both the discriminative capacity and generalization ability of the model across diverse scenarios.